OpenStudy (anonymous):

(x+2)log_55^x = X How do I solve this?

OpenStudy (anonymous):

\[(x+2)\log_{5}5^{x} = X \]

OpenStudy (anonymous):

Are you sure on right side you have capital X ?

OpenStudy (anonymous):

Same x on both sides

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

The power rule of logarithm says: \[\log(x)^y = y \cdot \log(x)\]

OpenStudy (anonymous):

\[(x+2)\log_{5}5^{x} = x \implies (x+2) \cdot x \cdot \log_5(5) = x\] Getting?

OpenStudy (anonymous):

bring it all to the front? why not to the exponent?

OpenStudy (anonymous):

but yes I get it

OpenStudy (anonymous):

Bring the exponent to the front in multiplication.. exponent is \(x\) there, so bring it front, okay?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The other rule says: If base and argument is same then value of logarithm is \(1\) ie: \[\log_a(a) = 1\]

OpenStudy (anonymous):

This way: \(\log_5(5) = ??\)

OpenStudy (anonymous):

so (x+2) * x * 1 = x?

OpenStudy (anonymous):

Yep.. :)

OpenStudy (anonymous):

so then x(x+2) = x?

OpenStudy (anonymous):

Yes, distribute x to brackets...

OpenStudy (anonymous):

x^2 + 2x = 2

OpenStudy (anonymous):

2 or x??

OpenStudy (anonymous):

\(x^2 + 2x = x\) Right?

OpenStudy (anonymous):

\[x^2 = -2x\]

OpenStudy (anonymous):

Wait, don't go fast..

OpenStudy (anonymous):

oops I did something wrong haha

OpenStudy (anonymous):

\(x^2 + 2x = x\) Subtract \(x\) from both the sides.. :)

OpenStudy (anonymous):

so x^2 + x = 0?

OpenStudy (anonymous):

Yep, take \(x\) common from left hand side now.

OpenStudy (anonymous):

x(x + 1) = 0

OpenStudy (anonymous):

Great, well done.. :)

OpenStudy (anonymous):

Now, can you find \(x\) ??

OpenStudy (anonymous):

See, when this situation happens : \((x+a)(x+b) = 0\) Then either (x+a) = 0 or (x+b) = 0.. Use this..

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

and x = 0

OpenStudy (anonymous):

Thats the answer in my book haha Thanks! :)

OpenStudy (anonymous):

So the two answers are : \(x =0,1\)

OpenStudy (anonymous):

\(\text{This is it}..\)